Managers of Crane Embroidery have decided to purchase a new monogram machine and are considering two alternative machines. The first machine costs $102,000 and is expected to last five years. The second machine costs $163,000 and is expected to last eight years. Assume that the opportunity cost of capital is 8 percent. What is the equivalent annual cost for each system?
Answer #1

Equivalent annual cost of Machine first = initial investment/PVAF at 8% for 5 years

= -102000/3.9927 =25546.62

Initial Investment = -102000

PVAF at 8% for 5 Years = 1-(1+r)^-n /r = 1-(1.08)^-5 / .08 =3.9927

Equivalent annual cost of Machine 2= initial investment/PVAF at 8% for 8 years

= -163000/5.7466 =25546.62

Initial Investment = -163000

PVAF at 8% for 5 Years = 1-(1+r)^-n /r = 1-(1.08)^-8 / .08 = 5.7466